OpenStudy (anonymous):

midpoint H(13,8),X(-6,-6) i got two different answers the fist time i did i got 3.5, 1 then got this answer 2 , 1 which one is right 7

OpenStudy (austinl):

\(\normalsize{Midpoint}=\Large{(\frac{x_1+x_2}{2},\frac{y_1+y_2}{2})}\)

OpenStudy (anonymous):

 so is that right

OpenStudy (anonymous):

3.5, 1 is right

OpenStudy (austinl):

Yes, but again exact forms are typically best.

OpenStudy (anonymous):

so what one is it

OpenStudy (austinl):

You have the correct answer, but you have it in "approximate" form.

OpenStudy (anonymous):

10.5 and -6 is it right

OpenStudy (austinl):

\(\normalsize{Midpoint}=\Large{(\frac{13+(-6)}{2},\frac{8+(-6)}{2})}\)

OpenStudy (anonymous):

(3.5, 1) i was right the first time wow lol

OpenStudy (austinl):

That is what I said -_-

OpenStudy (anonymous):

sorry